0.21 kg of gas is heated by means of an electric heater for a period of 10 mins during which time the pressure of the gas remains constant. the temperature of the gas is increased from 160C to 780C. the power used by the heater is 20 watts assuming no losses determine

a)the specific heat capacity of the gas at constant pressure

b)the specific heat capacity of the gas at constant volume

c)the characteristic gas constant

d)the density of the gas at a temperature of 160 C and with a pressure of o.12MN/m2


for the gas, take y=1.38



1
Expert's answer
2021-10-15T11:20:02-0400

"a)Q=P*t=mC_p\u2206T"

"C_p=\frac{Pt}{m∆T}"

"=\frac{20*10*60*10^{-3}}{0.21*(78-16)}"

=0.9217kJ/kgK

b)

"\gamma=\frac{C_p}{C_v}"

"C_v=\frac{C_p}{\gamma}"

="\frac{0.9217}{1.38}"

=0.6679kJ/kgK

c)

"R=C_p-C_v"

=0.9217-0.6679

=0.2538kJ/kgK

d)

"\rho=\frac{P}{RT}"

"=\frac{0.12*10^{6}}{0.2538*10^{3}*(16+273)}"

=1.6360kg/m3